# Fluctuations of Young diagrams for symplectic groups and semiclassical orthogonal polynomials

### v2

@@ 5. Conclusion and future work
In this paper, we have investigated the local fluctuations of random Young diagrams arising from the skew Howe duality for symplectic groups $Sp_{2n}\times Sp_{2k}$. The probability measure on diagrams, proportional to the ratio of the dimension of the irreducible $Sp_{2n}\times Sp_{2k}$-representation to the dimension of the exterior algebra $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)$, was expressed as a determinantal point process governed by a correlation kernel derived from semiclassical orthogonal polynomials. Unlike the $GL(n)$ case, where free-fermionic methods apply, the symplectic setting required a Christoffel transformation of Krawtchouk polynomials to construct the relevant orthogonal polynomials. It is an open question how to apply the free-fermionic representation [44] to the skew $Sp_{2n}\times Sp_{2k}$ case. In particular, it is not clear how the tau functions for d-connections can be constructed [45, 46].
@@
By analyzing the integral representation of these polynomials in the double-scaling limit $n,k\to\infty$ with $n/k$ fixed, we have demonstrated that the correlation kernel converges to the discrete sine kernel in the bulk regime. This universality mirrors results for general linear groups [11, 13] but required novel techniques due to the absence of free-fermionic structures.
Our findings align with the limit shape results in [22] and extend the understanding of fluctuations to symplectic symmetry classes. Future work could explore edge asymptotics and connections to discrete Painleve equations. The interplay between representation theory, determinantal processes in classical groups and the discrete integrable systems remains a rich area for further investigation.
6. Appendix
@@
The goal of this appendix is to prove the relation for the second term in the definition (2.9)
$\frac{\widetilde{K}_{m+2}(0)}{\widetilde{K}_{m}(0)}=-\frac{(K-m)(m+1)}{4n^{2}}=\frac{(m+1)(m+2)\binom{K}{m+2}\,_{2}F_{1}\left(-\frac{K}{2},-m-2;-K;2\right)}{4n^{2}\binom{K}{m}\,_{2}F_{1}\left(-\frac{K}{2},-m;-K;2\right)}.$ (6.1)
First of all we need to use the following transformation for $n\in\mathbb{N}$ (see (3.19) in [47])
@@
Substituting this into the original expression, we immediately come to the desired proof.

### v3

@@ 5 Conclusion and future work
In this paper, we have investigated the local fluctuations of random Young diagrams arising from the skew Howe duality for symplectic groups $Sp_{2n}\times Sp_{2k}$. The probability measure on diagrams, proportional to the ratio of the dimension of the irreducible $Sp_{2n}\times Sp_{2k}$-representation to the dimension of the exterior algebra $\bigwedge\left(\mathbb{C}^{2n}\otimes\mathbb{C}^{k}\right)$, was expressed as a determinantal point process governed by a correlation kernel derived from semiclassical orthogonal polynomials. Unlike the $GL_{n}$ case, where free-fermionic methods apply, the symplectic setting required a Christoffel transformation of Krawtchouk polynomials to construct the relevant orthogonal polynomials. It is an open question how to apply the free-fermionic representation [Baker_1996] to the skew $Sp_{2n}\times Sp_{2k}$ case. In particular, it is not clear how the tau functions for $d$-connections can be constructed [10.1215/S0012-7094-06-13433-6, arinkin2009tau].
@@
By analyzing the integral representation of these polynomials in the double-scaling limit $n,k\to\infty$ with $n/k$ fixed, we have demonstrated that the correlation kernel converges to the discrete sine kernel in the bulk regime. This universality mirrors results for general linear groups [Borodin_2003, Borodin_2017] but required novel techniques due to the absence of free-fermionic structures.
Our findings align with the limit shape results in [nazarov2024skew] and extend the understanding of fluctuations to symplectic symmetry classes.
The following open problems and directions naturally follow from the present work:
  1. 1.
    Free-fermionic representation for the skew $Sp_{2n}\times Sp_{2k}$ case. Clarify whether and how the free-fermionic formalism used in the $GL_{n}$ setting can be adapted to the symplectic skew Howe duality. See [Baker_1996] for related representations.
  2. 2.
    Tau functions for $d$-connections. Construct the appropriate tau functions in the symplectic skew-duality setup and explain their relation to the determinantal kernel obtained here [10.1215/S0012-7094-06-13433-6, arinkin2009tau].
  3. 3.
    Edge asymptotics and connections to discrete Painlevé equations. Study edge scaling regimes (soft and hard edges) to determine whether discrete analogues of Airy/Tracy–Widom laws or Painlevé-type distributions arise, and relate these asymptotics to discrete Painlevé equations where applicable.
6 Appendix
@@
The goal of this appendix is to prove the relation for the second term in the definition (2.9).
Lemma 4.
Polynomials $\widetilde{K}_{m}(x)$, defined by (2.4), satisfy the following relation:
$\frac{\widetilde{K}_{m+2}(0)}{\widetilde{K}_{m}(0)}$ $=-\frac{(K-m)(m+1)}{4n^{2}}=$
$=\frac{(m+1)(m+2)\binom{K}{m+2}\,_{2}F_{1}\left(-\frac{K}{2},-m-2;-K;2\right)}{4n^{2}\binom{K}{m}\,_{2}F_{1}\left(-\frac{K}{2},-m;-K;2\right)}.$ (6.1)
Proof.
First of all we need to use the following transformation for $n\in\mathbb{N}$ (see (3.19) in [Hannah2013IdentitiesFT])
@@
Substituting this into the original expression, we immediately come to the desired proof. ∎
Aknowlegments We thank Travis Scrimshaw for reading the first draft and pointing out the deficiencies in it. We thank Anton Dzhamay for useful discussions.
Funding The study was carried out with the financial support of the Ministry of Science and Higher Education of the Russian Federation in the framework of a scientific project under agreement No. 075-15-2025-013.
Data Availability All data generated or analysed during this study are contained in this document.
Statements and Declarations
Conflict of interest On behalf of all authors, the corresponding author states that there is no conflict of interest.